OpenStudy (anonymous):

what is 1.3 repeating as a fraction

OpenStudy (ahsome):

Here is one way: \[\frac{13}{10}\]

OpenStudy (anonymous):

thanks :)

OpenStudy (ahsome):

No problem :)

OpenStudy (anonymous):

4/3 actually. 13/10=1.3 exactly.

OpenStudy (anonymous):

im confused now

OpenStudy (anonymous):

4/3=1.3 repeating 13/10=1.3

OpenStudy (ahsome):

Isn't that what he want's. I guessed he wanted 1.3, not 1.33333333333333333333333333333333333333333333333333333etc

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

It said 1.3 repeating.

OpenStudy (ahsome):

Oh, whoops. Then: \[\frac{1}{3}\] Or any multiple of that: \[\frac{2}{6}\] \[\frac{3}{9}\] ETC

OpenStudy (anonymous):

thanks both of you :)

OpenStudy (ahsome):

No Problem :)